OpenStudy (anonymous):

Differentiate: ln(x-1/x^3+1)

OpenStudy (anonymous):

I really need help with this one! natural logs always throw me off.

OpenStudy (anonymous):

so I know I can split this equation because of the natural logs rule like this: \[(lnx-1)-(lnx ^{3}+1)\]

OpenStudy (anonymous):

but then do i take the derivative of each term? which rule do I use?

zepdrix (zepdrix):

\[\Large \ln(x-1)-\ln(x^3+1)\]Yah looks good so far :) Hmm, yes we want to take the derivative of each term separately. We'll use the rule for the derivative of natural log :O remember that one?

OpenStudy (anonymous):

which would be inverse right?

OpenStudy (anonymous):

so I would have something like 1/(x-1)

zepdrix (zepdrix):

Gives us the reciprocal thing? yah that looks good for the first one. Don't forget to apply the `chain rule` on the second term.

OpenStudy (anonymous):

after I do the reciprocal thing for the second term right?

OpenStudy (anonymous):

also how do I know I'm suppose to do the chain rule for the second only?

zepdrix (zepdrix):

You can apply the chain rule to the first term also, it might be good practice.\[\Large \left[\ln(x-1)\right]' \qquad=\qquad \frac{1}{x-1}\color{royalblue}{(x-1)'} \qquad=\qquad \frac{1}{x-1}\color{royalblue}{(1)}\]

zepdrix (zepdrix):

You want to try and think ahead, "If I apply the chain rule, will it give me anything more than just 1?" If the answer is yes, then you need to apply the chain rule.

zepdrix (zepdrix):

See how applying it to the first term gave us 1 for the derivative of the inner function? :o

OpenStudy (anonymous):

so we didnt' need to use the chain rule

zepdrix (zepdrix):

Didn't need to, correct ^^

OpenStudy (anonymous):

ok so the second term needs to use the chain rule right?

zepdrix (zepdrix):

Yes, because the derivative of the inner function will end up giving us more than just 1.

OpenStudy (anonymous):

right. can u check if I have the right answer?

zepdrix (zepdrix):

sure c: whatcha got?

OpenStudy (anonymous):

for the second term I got 3x/x^3+1

zepdrix (zepdrix):

Woops, small mistake with your chain rule there! :) \[\large \left[\ln(x^3+1)\right]' \qquad=\qquad \frac{1}{x^3+1}\color{royalblue}{(x^3+1)'} \qquad=\qquad \frac{1}{x^3+1}\color{royalblue}{(3x^2)}\]

OpenStudy (anonymous):

oh yeah I see it :) haha

OpenStudy (anonymous):

so then is this my final answer?

zepdrix (zepdrix):

Yes. You could combine the fractions into one big messy thing. But given that the problem worked out the way it did, with the logs splitting up, this is probably where you're expected to stop.

OpenStudy (anonymous):

YAY! Thanks!

OpenStudy (anonymous):

Oh I have one quick question

OpenStudy (anonymous):

its not about natural logs..

zepdrix (zepdrix):

Uh oh +_+ what kind of logs is it about then? I'm not a beaver.. i don't know much about wood besides the natural log D:

OpenStudy (anonymous):

oh my... lol its implicit differentiation problem

zepdrix (zepdrix):

Oh those are fun :)

OpenStudy (anonymous):

so I have y-x=sinxy

OpenStudy (anonymous):

yes haha sometimes?

zepdrix (zepdrix):

Like this? both x and y inside the sine? \[\Large y-x=\sin(xy)\]

OpenStudy (anonymous):

I'm wondering when I differentiate each term for sinxy do I just (cosxy)*y'?

OpenStudy (anonymous):

yup!

zepdrix (zepdrix):

It's gonna be a smidge more complex than that I'm afraid :u

zepdrix (zepdrix):

\[\Large \left[\sin(xy)\right]' \qquad=\qquad \left[\cos(xy)\right]\color{royalblue}{(xy)'}\]Here is our chain rule setup. Looks like we'll have to apply the product rule on this blue term.

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

so for the blue term its y+x?

OpenStudy (anonymous):

what do I do with y'? I just assumed that became 1.

zepdrix (zepdrix):

\[\Large =\cos(xy)\left[\color{royalblue}{(x)'}y+x\color{royalblue}{(y)'}\right]\] So you've got the right idea with your product rule setup. But noooo, y' doesn't become 1! :O That is the thing we're trying to solve for. x' becomes 1, because we're taking our derivative `with respect to x`. y' is the derivative term we're looking for. \[\Large =\cos(xy)\left[\color{royalblue}{(1)}y+x\color{royalblue}{y'}\right]\]

OpenStudy (anonymous):

ok .. so its just y+xy

OpenStudy (anonymous):

xy' i mean

OpenStudy (anonymous):

now what?

zepdrix (zepdrix):

\[\Large y-x=\sin(xy)\]Taking the derivative gives us,\[\Large y'-1=\cos(xy)\left(y+xy'\right)\]We want to distribute the cosine to each term in the brackets.

OpenStudy (anonymous):

ok~???

zepdrix (zepdrix):

\[\Large y'-1=y\cos(xy)+x \cos(xy)y'\] Understand that step ok? I know it looks a little messy.

zepdrix (zepdrix):

We need to get our (y')'s onto one side and this is the only way we can do it :o

OpenStudy (anonymous):

yeah I got it! but how are we gonna get our y's to one side..

zepdrix (zepdrix):

Let's subtract the (big ugly cosine with the y' attached to it) from each side.\[\Large y'-x \cos(xy)y'-1=y \cos(xy)\]Then we'll add 1 to each side,\[\Large y'-x \cos(xy)y'=y \cos(xy)+1\]Then we'll factor y' out of each term on the left,\[\Large y'\left[1-x \cos(xy)\right]=y \cos(xy)+1\] Understand that last step there? That one is pretty important. We can factor derivative terms just like any other variable.

OpenStudy (anonymous):

yup i'm following!

zepdrix (zepdrix):

to solve for y' we simply divide by that big mess on the left,\[\Large y'=\frac{y \cos(xy)+1}{1-x \cos(xy)}\]

OpenStudy (anonymous):

so ur just trying first to get y' on each side not just any y's.

zepdrix (zepdrix):

Correct just `y'` 's, we don't care about the y's :)

zepdrix (zepdrix):

Stupid apostrophe :P man that's confusing lolol

OpenStudy (anonymous):

yup got it!

OpenStudy (anonymous):

wow thanks for the help! ok one last question lol this one you don't have to walk through the whole thing with me cuz you've been such a big help but i was just wondering...

zepdrix (zepdrix):

heh

OpenStudy (anonymous):

when you have three terms like this (xsomething)(xsomething)(ysomething) do you just do the product rule? Implicit differentiation again.

zepdrix (zepdrix):

Your text book might suggest that you apply logarithmic differentiation. Because:\[\large \ln(x^3e^{2x}\cos x) \qquad=\qquad \ln x^3+ \ln(e^{2x})+\ln(\cos x)\] That would allow you to split up terms easily and take their derivatives individually. But yah another option is to apply the product rule. It follows the same pattern with 3 terms.\[\Large (uvw)' \qquad=\qquad u'vw+uv'w+uvw'\]

zepdrix (zepdrix):

Oh sorry I misread your comment :3 thought the last one was (xsomething). No big deal though, same idea applies. Probably would want to do product rule if it involves (ysomething).

OpenStudy (anonymous):

ok thanks so much for ur help!!